# Singularity of the axisymmetric stagnation-point-like solution within a cylinder of the 3D Euler incompressible fluid equations

### v1

@@ 2.2 Solutions for Plane Vorticity, Stretching Rate, Pathlines, and Velocity Field
Using the area-preserving identity:
@@
we derive the relation:
$\left\langle J\frac{r}{r_{0}}\right\rangle_{0}=1,$ (25)
where the angular brackets are defined in equation (19).
Differentiating the area identity with respect to time yields the evolution equation for the Jacobian:
$\frac{\dot{J}}{J}=\frac{\partial u_{r}}{\partial r},\quad J(0)=1,$ (26)
which is a first-order linear ordinary differential equation. Combining this with the expression for $J$ in equation (23), we obtain the following partial differential equation for $r(r_{0},t)$:
Solving this equation yields the radial pathline function $r(r_{0},t)$, consistent with the expression stated in Theorem 1. Furthermore, by combining equations (25) and (27), we recover the ordinary differential equation governing $S(t)$, which was introduced in equation (18).
Additionally, using the relation $u_{\theta}=r\frac{\partial\theta}{\partial t}$ from equation (9) and the Lagrangian solution for the plane vorticity $\omega$ given in equation (17), we deduce the angular component of the pathline $\theta(r_{0},\theta_{0},t)$, matching the expression provided in Theorem 1.
@@
The Eulerian solution can be obtained by expressing the initial position $\bm{X}_{0}$ as a function of the current position $\bm{X}$ and time $t$, i.e., $\bm{X}_{0}=\bm{X}_{0}(\bm{X},t)$, and substituting this into the Lagrangian solution. However, for the ASCII fluid, the mapping $\bm{X}_{0}=\bm{X}_{0}(\bm{X},t)$ is implicit due to the structure of the pathline equations (21). As a result, the Eulerian solution cannot generally be expressed in an explicit closed form. Nevertheless, under certain specific Lagrangian solutions, it is possible to derive explicit Eulerian expressions, one of which will be presented in section 3.1.
@@
From equation (18), we observe that the temporal evolution of the function $S(t)$ depends solely on the initial stretching rate $\gamma_{0}(r_{0})$. Consequently, Theorem 1 implies that each pair of Lagrangian vorticity and stretching rate $(\gamma,\omega)$ uniquely determines a corresponding horizontal pathline $(r,\theta)$, a property analogous to that of the Eulerian description. For clarity, we define the pair $(\gamma,\omega)$ as the vorticity coordinate. Based on this observation, once the vorticity coordinate is determined, the full pathline solution for the ASCII fluid becomes known. Therefore, our subsequent analysis will focus exclusively on the dynamics of the vorticity coordinate.
@@
There are two notable special cases in the pathline solution (21) of the ASCII fluid:
@@
    Pathlines starting at the centre: If the initial radial position is $r_{0}=0$, then the radial component of the pathline satisfies:
@@
3 Singularity and Asymptotic Behavior of ASCII Fluid Fields
@@
The singularity time $T_{*}$ can be either finite or infinite. An infinite value of $T_{*}$ implies the absence of a finite-time singularity. Since $S(t)\geq 0$ for all $t\geq 0$, the singularity condition $1+\gamma_{0}(r_{0})S(t)=0$ is first satisfied at the minimum point of the initial stretching rate $\gamma_{0}(r_{0})$. Therefore, we conclude that any blowup must occur at the minimum point(s) of $\gamma_{0}(r_{0})$.
This leads us to define the blowup pathline as the fluid trajectory starting from the minimum point $r_{0}=r_{-}:=\mathrm{argmin}(\gamma_{0})$. In other words, the location of the singularity is determined entirely by the structure of the initial stretching rate $\gamma_{0}(r_{0})$, specifically its minimum value.
@@
Using the evolution equation for $S(t)$ given in equation (18), we derive an expression for the singularity time:
$T_{*}=\int_{0}^{S_{*}}\frac{1}{\dot{S}}\,dS=\int_{0}^{S_{*}}\left\langle\frac{1}{1+\gamma_{0}(r_{0})S}\right\rangle_{0}^{2}dS,$ (34)
@@
It is important to emphasise that our analysis assumes regular initial conditions. We begin by examining the case where $\gamma_{0}(r_{0})$ takes the form of an exact parabola, and then proceed to more general cases where $\gamma_{0}(r_{0})$ is continuous but not necessarily differentiable near the blowup point.
@@ 3.1 Parabolic Initial Stretching Rate $\gamma_{0}(r_{0})$
    The minimum and supremum of $\gamma_{0}$ are negatives of each other, leading to the same value for the minimum: $\inf\gamma_{0}=-|{f}|$, regardless of whether f{f} is positive or negative;
@@
Firstly, substituting this form of $\gamma_{0}(r_{0})$ into the ODE governing $S(t)$ given in equation (18), we obtain:
$\dot{S}(t)=4{f}^{2}S(t)^{2}\left[\ln\left|\frac{1-{f}S(t)}{1+{f}S(t)}\right|\right]^{-2}.$ (37)
It is important to note that changing the sign of f{f} does not alter the solution of $S(t)$, as the right-hand side of the above equation remains invariant under such a transformation. Furthermore, from equation (34) we derive an explicit expression for the singularity time:
@@
$\gamma(r_{0}=r_{-},t)\sim-\frac{1}{T_{*}-t},\qquad\omega(r_{0}=r_{-},t)$ $\sim$ $\omega_{0}(r_{-})\sqrt{|f|(T_{*}-t)}\,\mathrm{e}^{W_{-1}\left(-\sqrt{|f|(T_{*}-t)}\right)},$
@@
namely $S_{*}=1/|{f}|$. Using this relation together with equations (37) and (20), we derive the leading-order asymptotic expressions for the first and second time derivatives of $S(t)$:
@@
The explicit asymptotic solution for $S(t)$ near the singularity time $T_{*}$ can be obtained from the integral expression:
$T_{*}-t=\int_{S}^{S_{*}}\frac{1}{\dot{S}}\mathrm{d}S,$ (46)
which yields the leading-order approximation:
@@
Substituting this into the Lagrangian expressions for $\gamma$ and $\omega$ given in equation (17), we compute the asymptotic behavior of the stretching rate along the blowup pathline:
$\gamma(r_{0}=r_{-},t)=\frac{\gamma_{0}\dot{S}}{1+\gamma_{0}S}-\frac{\ddot{S}}{2\dot{S}}\sim-\frac{4\left[\ln\left(1-\frac{S}{S_{*}}\right)\right]^{-2}}{S_{*}-S}-\frac{4\left[\ln\left(1-\frac{S}{S_{*}}\right)\right]^{-3}}{S_{*}-S}\sim-\frac{1}{T_{*}-t}.$ (48)
Similarly, for the plane vorticity we obtain:
$\omega(r_{0}=r_{-},t)\sim\frac{\omega_{0}(r_{-})(1+\gamma_{0}S)}{\sqrt{\dot{S}}}\sim\frac{\omega_{0}(r_{-})\left(1-\frac{S}{S_{*}}\right)}{2\left|\ln\left(1-\frac{S}{S_{*}}\right)\right|^{-1}}\sim\omega_{0}(r_{-})\sqrt{|{f}|(T_{*}-t)}\,\mathrm{e}^{W_{-1}\left(-\sqrt{|{f}|(T_{*}-t)}\right)}.$ (49)
@@
As noted by Gibbon and Ohkitani, a necessary condition for the breakdown of the Gibbon-type model—similar in spirit to the BKM theorem—is that the time integral of the $L^{\infty}$-norm of the stretching rate diverges:
@@
This criterion is exemplified by Theorem 2, which demonstrates that such a divergence occurs precisely at the blowup time $T_{*}$ (Beale et al., 1984; Ohkitani & Gibbon, 2000).
@@
    Case f<0{f}<0: Blowup at the centre $r=0$: When f<0{f}<0, the minimum point of the initial stretching rate is located at $r_{-}=0$. In this case, we have:
@@
  • $\gamma(r,t)$ $\sim-\frac{1}{T_{*}-t}e^{\frac{2r^{2}}{R^{2}}W_{-1}\left(-\sqrt{|f|(T_{*}-t)}\right)}-\frac{1}{2(T_{*}-t)\left|W_{-1}\left(-\sqrt{|f|(T_{*}-t)}\right)\right|},$ (52)
@@
    From the expression for $\gamma(r,t)$, we observe that the stretching rate blows up over the entire spatial domain $\Omega_{C}$, with the most rapid growth occurring at the origin $r=0$, where $\gamma(0,t)\sim-\frac{1}{T_{*}-t}$, confirming that the singularity first emerges at the centre. An example of this asymptotic behavior and its evolution for the case f=10,R=1{f}=-10,R=1 is illustrated in Figure 1.
    In contrast, the plane vorticity $\omega(r,t)$ behaves differently: it decays to zero throughout the interior of the domain but diverges at the boundary $r=R$. This boundary blowup occurs more slowly than the blowup of stretching rate at the centre, with a scaling that is proportional to the negative branch of the Lambert $W$ function: $W_{-1}\left(-\sqrt{|{f}|(T_{*}-t)}\right)$. Despite the slower growth rate, the presence of a finite-time singularity at the boundary confirms the non-regularity of the solution;
@@
    Similar to the previous case, here the stretching rate $\gamma(r,t)$ becomes singular everywhere on the spatial domain $\Omega_{C}=[0,R]\times[0,2\pi)\times\mathbb{R}$ at $t=T_{*}$, but with the fastest growth rate occurring at the boundary $r=R$, where $\gamma(R,t)\sim-\frac{1}{2(T_{*}-t)}$. This behavior is demonstrated numerically in Figure 2 for f=10{f}=10, $R=1$.
    As for the plane vorticity $\omega(r,t)$, it decays to zero at the singularity time everywhere except at the centre of the domain $r=0$, where it blows up. This indicates that the singularity formation mechanism differs depending on the sign of the initial stretching rate f{f}, even though both cases share the same minimum value and singularity time.
@@
Figure 1: Evolution of the parabolic stretching rate $\gamma_{0}(r_{0})=20r_{0}^{2}-10$ (hence f=10,R=1f=-10,R=1) in the Eulerian description. The singularity time is computed as $T_{*}=\frac{\pi^{2}}{6|{f}|}\approx 0.16449$. According to equation (52), the asymptotic behavior of the stretching rate $\gamma$ is expected to take the form of a bell-shaped function, which is consistent with the spatial profile observed in this figure.
These results demonstrate that the parabolic initial profile of the stretching rate can lead to finite-time singularities in both Lagrangian and Eulerian descriptions of the ASCII fluid. The location and nature of the blowup are highly sensitive to the sign of f{f}, despite the fact that the minimum of $\gamma_{0}(r_{0})$ remains unchanged under the transformation $\gamma_{0}\mapsto-\gamma_{0}$. One way to understand physically this difference in behaviour is to notice that when $f<0$ the blowup occurs at a point ($r_{-}=0$), whereas when $f>0$ the blowup occurs on a curve (the ring $r_{-}=R$).
@@
Figure 2: Evolution of the parabolic stretching rate $\gamma_{0}(r_{0})=10-20r_{0}^{2}$ (hence $f=10,R=1$) in the Eulerian description. According to equation (54), the asymptotic behavior of the stretching rate $\gamma$ is expected to exhibit a highly skewed profile near the boundary as the solution approaches the singularity time. This characteristic is clearly reflected in the spatial distribution observed in the figure.

### v2

@@ 2.2 Solutions for Plane Vorticity, Stretching Rate, Pathlines, and Velocity Field
The evolution equation for the Jacobian determinant follows from its definition as the rate of change of area under the flow map. Specifically, differentiating $J=\frac{\partial r}{\partial r_{0}}$ with respect to time and using the radial pathline evolution equation $\mathrm{d}r/\mathrm{d}t=u_{r}(r,t)$ yields
$\frac{\dot{J}}{J}=\frac{\partial u_{r}}{\partial r},\quad J(0)=1,$ (25)
which is a first-order linear ordinary differential equation.
To derive the radial pathline solution $r(r_{0},t)$, we note that the LHS of the Jacobian evolution equation (25) is a material time derivative. It turns out that the RHS of that equation is also a material time derivative. This will allow us to find a first integral. To see how the RHS is a material time derivative, we use the incompressibility condition (8) to get $\frac{\partial u_{r}}{\partial r}=-\gamma-\frac{u_{r}}{r}$. Now, from the Lagrangian solution (18) it is clear that $\gamma$ is a material time derivative. Finally, from the pathline definition $\dot{r}=u_{r}(r,t)$, we get $\frac{u_{r}}{r}=\frac{\dot{r}}{r}$, again a material time derivative. Substituting these relations, integrating with respect to time from 0 to $t$, and then exponentiating, yields the following ordinary differential equation for the radial map:
Solving this equation yields the radial pathline function $r(r_{0},t)$, consistent with the expression (22) stated in Theorem 1. Furthermore, using the area-preserving identity:
@@
and replacing $J$ from equation (24), we derive the relation:
$\left\langle\frac{r}{r_{0}}\frac{\partial r}{\partial r_{0}}\right\rangle_{0}=1,$ (28)
where the angular brackets are defined in equation (20). Combining equations (28) and (26), we obtain the ordinary differential equation governing $S(t)$, which was introduced in equation (19).
Next, we combine the angular pathline evolution equation $u_{\theta}=rd\theta/dt$ with the definition (9) of $u_{\theta}$, the result (26) for the Jacobian, and the Lagrangian solution (18) for $\omega$, to obtain the angular pathline function $\theta(r_{0},\theta_{0},t)$ via quadrature, matching the expression (22) stated in Theorem 1.
@@
The Eulerian solution can be obtained by expressing the initial position $\bm{X}_{0}$ as a function of the current position $\bm{X}$ and time $t$, i.e., $\bm{X}_{0}=\bm{X}_{0}(\bm{X},t)$, and substituting this into the Lagrangian solution. However, for the cylindrical Gibbon flow, the mapping $\bm{X}_{0}=\bm{X}_{0}(\bm{X},t)$ is implicit due to the structure of the pathline equations (22). As a result, the Eulerian solution cannot generally be expressed in an explicit closed form. Nevertheless, under certain specific Lagrangian solutions, it is possible to derive explicit Eulerian expressions, one of which will be presented in section 3.1.
@@
From equation (19), we observe that the temporal evolution of the function $S(t)$ depends solely on the initial stretching rate $\gamma_{0}(r_{0})$. Consequently, Theorem 1 implies that each pair of Lagrangian vorticity and stretching rate $(\gamma,\omega)$ uniquely determines a corresponding horizontal pathline $(r,\theta)$, a property analogous to that of the Eulerian description. For clarity, we define the pair $(\gamma,\omega)$ as the vorticity coordinate. Based on this observation, once the vorticity coordinate is determined, the full pathline solution for the cylindrical Gibbon flow becomes known. Therefore, our subsequent analysis will focus exclusively on the dynamics of the vorticity coordinate.
@@
There are two notable special cases in the pathline solution (22) of the cylindrical Gibbon flow:
@@
    Pathlines starting at the center: If the initial radial position is $r_{0}=0$, then the radial component of the pathline satisfies:
@@
3 Singularity and Asymptotic Behavior of Cylindrical Gibbon Flow
@@
Remark 3.0.1
In principle, the condition (32) does not necessarily imply a blowup of the stretching rate $\gamma$, because there exists a special scenario in which the two terms $\frac{\gamma_{0}(r_{0})\dot{S}(t)}{1+\gamma_{0}(r_{0})S(t)}$ and $\frac{\ddot{S}(t)}{2\dot{S}(t)}$ diverge at exactly the same rate, leading to their cancellation and leaving $\gamma$ regular, see the expression of $\gamma$ in (18). However, such a cancellation does not occur in the analysis presented below; hence, within the context of this paper, the fulfillment of (32) indeed implies a genuine blowup of $\gamma$.
The singularity time $T_{*}$ can be either finite or infinite. An infinite value of $T_{*}$ implies the absence of a finite-time singularity. At each time $t>0$, the fact $S(t)>0$ implies: $1+\gamma_{0}(r_{0})S(t)<1+\gamma_{0}(r_{1})S(t)$ if and only if $\gamma_{0}(r_{0})<\gamma_{0}(r_{1})$. Now, because $S(0)=0$ the functions of time $h(r_{0},t):=1+\gamma_{0}(r_{0})S(t)$ all start at the value $1$ at $t=0$: $h(r_{0},0)=1$. These functions are ordered by the previous result, namely: when $t>0$ we have $h(r_{0},t)<h(r_{1},t)$ if and only if $\gamma_{0}(r_{0})<\gamma_{0}(r_{1})$. As these functions are continuous functions of time $t$ (because $S$ is continuous), the first function that can take the value 0 is the one for which $\gamma_{0}(r_{0})$ is minimal, and we will have $1+\min_{r_{0}\in[0,R]}\gamma_{0}(r_{0})S(T_{*})=0$, with $T_{*}$ the time at which this happens. As $S(T_{*})>0$, a necessary condition for a finite-time singularity is $\min_{r_{0}\in[0,R]}\gamma_{0}(r_{0})<0$, which is satisfied in our case (except in the trivial case $\gamma_{0}\equiv 0$) because $\langle\gamma_{0}\rangle_{0}=0$. We denote a point where $\gamma_{0}(r_{0})$ attains its minimum value as $r_{-}:=\arg\min_{r_{0}\in[0,R]}\gamma_{0}(r_{0})$. We conclude that any blowup must occur at the pathline starting at such a point $r_{-}$, where $\gamma_{0}(r_{0})$ attains its minimum value. This leads us to define the blowup pathline as the fluid trajectory starting from the minimum point $r_{0}=r_{-}$.
@@
Integrating the evolution equation (19) for $S(t)$ by separation of variables yields an explicit expression for the singularity time:
$T_{*}=\int_{0}^{S_{*}}\frac{1}{\dot{S}}\,\mathrm{d}S=\int_{0}^{S_{*}}\left\langle\frac{1}{1+\gamma_{0}(r_{0})S}\right\rangle_{0}^{2}\mathrm{d}S,$ (35)
@@
It is important to emphasize that our analysis assumes regular initial conditions. We begin by examining the case where $\gamma_{0}(r_{0})$ takes the form of an exact parabola, and then proceed to more general cases where $\gamma_{0}(r_{0})$ is continuous but not necessarily differentiable near the blowup point.
@@ 3.1 Parabolic Initial Stretching Rate $\gamma_{0}(r_{0})$
    The minimum and supremum of $\gamma_{0}$ are negatives of each other, leading to the same value for the minimum: $\min\gamma_{0}=-|{f}|$, regardless of whether f{f} is positive or negative;
@@
Firstly, substituting this form of $\gamma_{0}(r_{0})$ into the expressions for $\dot{S}(t)$ and $\ddot{S}(t)$ given in equations (19) and (21), we obtain:
$\dot{S}(t)$ $=4{f}^{2}S(t)^{2}\left[\ln\left|\frac{1-{f}S(t)}{1+{f}S(t)}\right|\right]^{-2}\,$, (38)
$\ddot{S}(t)$ $={32f^{4}S(t)^{3}\left[\ln\left|\frac{1-{f}S(t)}{1+{f}S(t)}\right|\right]^{-5}\left[\ln\left|\frac{1-{f}S(t)}{1+{f}S(t)}\right|+\frac{2fS(t)}{1-f^{2}S(t)^{2}}\right]}\,.$ (39)
It is important to note that changing the sign of f{f} does not alter the solution of $S(t)$, as the right-hand side of the above equation remains invariant under such a transformation. Furthermore, substituting formula (38) into the singularity time formula (35) and evaluating the resulting integral yields an explicit expression for the blowup time:
@@
$\gamma(r_{0}=r_{-},t)\sim-\frac{1}{T_{*}-t},\qquad\omega(r_{0}=r_{-},t)$ $\sim$ $\omega_{0}(r_{-}){\frac{|f|(T_{*}-t)}{\left|W_{-1}\left(-\sqrt{|f|(T_{*}-t)}\right)\right|},}$
@@
namely $S_{*}=1/|{f}|$. Substituting this relation into equations (38) and (39), we derive the leading-order asymptotic expressions for the first and second time derivatives of $S(t)$:
@@
An explicit asymptotic solution for $S(t)$ at times $t$ near the singularity time $T_{*}$ can be obtained via first replacing $\dot{S}$ into the formula for the singularity time:
$T_{*}-t$ $=\int_{S}^{S_{*}}\frac{1}{\dot{S}}\mathrm{d}S=\frac{1}{4}\int_{S}^{S_{*}}\left[\ln\left(1-\frac{S}{S_{*}}\right)\right]^{2}dS\sim\frac{1}{4}(S_{*}-S)\left[\ln\left(S_{*}-S\right)\right]^{2}$ (48)
$=\left(\sqrt{S_{*}-S}\ln\sqrt{S_{*}-S}\right)^{2}\,$,
and then applying the definition of the negative branch of the Lambert $W$ function as in the statement of this theorem, yields the leading-order approximation
@@
Substituting the asymptotics (47) and (49) into the Lagrangian expressions for $\gamma$ and $\omega$ given in equation (18), we compute the asymptotic behavior of the stretching rate along the blowup pathline:
$\gamma(r_{0}=r_{-},t)=\frac{\gamma_{0}\dot{S}}{1+\gamma_{0}S}-\frac{\ddot{S}}{2\dot{S}}\sim-\frac{4\left[\ln\left(1-\frac{S}{S_{*}}\right)\right]^{-2}}{S_{*}-S}-\frac{4\left[\ln\left(1-\frac{S}{S_{*}}\right)\right]^{-3}}{S_{*}-S}\sim-\frac{1}{T_{*}-t}\,$, (50)
and for the plane vorticity we obtain:
$\omega(r_{0}=r_{-},t)\sim\frac{\omega_{0}(r_{-})(1+\gamma_{0}S)}{\sqrt{\dot{S}}}\sim\frac{\omega_{0}(r_{-})\left(1-\frac{S}{S_{*}}\right)}{2\left|\ln\left(1-\frac{S}{S_{*}}\right)\right|^{-1}}\sim\omega_{0}(r_{-}){\frac{|f|(T_{*}-t)}{\left|W_{-1}\left(-\sqrt{|f|(T_{*}-t)}\right)\right|}.}$ (51)
@@
As noted by Gibbon and Ohkitani, a necessary condition for the breakdown of the Gibbon-type model—similar in spirit to the Beale-Kato-Majda theorem—is that the time integral of the $L^{\infty}$-norm of the stretching rate diverges:
@@
This criterion is exemplified by Theorem 2, which demonstrates that such a divergence occurs precisely at the blowup time $T_{*}$ (Beale et al., 1984; Ohkitani and Gibbon, 2000).
@@
    Case f<0{f}<0: Blowup at the center $r=0$: When f<0{f}<0, the minimum point of the initial stretching rate is located at $r_{-}=0$. In this case, we have:
@@
    $\gamma(r,t)$ $\sim-\frac{1}{T_{*}-t}\mathrm{e}^{-\frac{2r^{2}}{R^{2}}\left|W_{-1}\left(-\sqrt{|f|(T_{*}-t)}\right)\right|}-\frac{1}{2(T_{*}-t)\left|W_{-1}\left(-\sqrt{|f|(T_{*}-t)}\right)\right|},$ (54)
@@
    From the expression for $\gamma(r,t)$, we observe that the stretching rate blows up over the entire spatial domain $\Omega_{C}$, with the most rapid growth occurring at the origin $r=0$, where $\gamma(0,t)\sim-\frac{1}{T_{*}-t}$, confirming that the singularity first emerges at the center.
    In contrast, the plane vorticity $\omega(r,t)$ behaves differently: it decays to zero throughout the interior of the domain but diverges at the boundary $r=R$. This boundary blowup occurs more slowly than the blowup of stretching rate at the center, with a scaling that is proportional to the negative branch of the Lambert $W$ function: $W_{-1}\left(-\sqrt{|{f}|(T_{*}-t)}\right)$. Despite the slower growth rate, the presence of a finite-time singularity at the boundary confirms the non-regularity of the solution;
    An example of the above asymptotic behavior for $f<0$ and its evolution is illustrated in Figure 1 in the case f=10,R=1{f}=-10,R=1.
@@
    Similar to the previous case, here the stretching rate $\gamma(r,t)$ becomes singular everywhere on the spatial domain $\Omega_{C}=[0,R]\times[0,2\pi)\times\mathbb{R}$ at $t=T_{*}$, but with the fastest growth rate occurring at the boundary $r=R$, where $\gamma(R,t)\sim-\frac{1}{2(T_{*}-t)}$.
    As for the plane vorticity $\omega(r,t)$, it decays to zero at the singularity time everywhere except at the center of the domain $r=0$, where it blows up. This indicates that the singularity formation mechanism differs depending on the sign of the initial stretching rate f{f}, even though both cases share the same minimum value and singularity time.
This behavior for $f>0$ is demonstrated numerically in Figure 2 for f=10{f}=10, $R=1$.
@@
$t=0$
$t=0.0822$
$t=0.1644$
Figure 1: Evolution of the parabolic initial stretching rate $\gamma_{0}(r_{0})=20r_{0}^{2}-10$ (hence f=10,R=1f=-10,R=1) and the uniform initial plane vorticity $\omega_{0}(r_{0})=10$ in the Eulerian description. The singularity time is computed as $T_{*}=\frac{\pi^{2}}{6|{f}|}\approx 0.16449$. According to equation (54), the asymptotic behavior of the stretching rate $\gamma$ is expected to take the form of a bell-shaped function, which is consistent with the spatial profile observed in this figure. While, at the center ($r=0$), $\gamma$ develops a finite-time singularity and $\omega$ develops a zero, at the boundary ($r=R$) $\omega$ exhibits a divergent peak, confirming their mutually exclusive blowup locations. The slower growth of $\omega$ compared to $\gamma$ (visible in the $z$-axis scale) aligns with its Lambert-W-type asymptotics derived in (54).
These results demonstrate that the parabolic initial profile of the stretching rate can lead to finite-time singularities in both Lagrangian and Eulerian descriptions of the cylindrical Gibbon flow. The location and nature of the blowup are highly sensitive to the sign of f{f}, despite the fact that the minimum of $\gamma_{0}(r_{0})$ remains unchanged under the transformation $\gamma_{0}\mapsto-\gamma_{0}$. One way to understand physically this difference in behavior is to notice that when $f<0$ the blowup occurs at a point ($r_{-}=0$), whereas when $f>0$ the blowup occurs on a curve (the ring $r_{-}=R$).
@@
$t=0$
$t=0.0822$
$t=0.1644$
Figure 2: Evolution of the parabolic stretching rate $\gamma_{0}(r_{0})=10-20r_{0}^{2}$ (hence $f=10,R=1$) and the uniform initial plane vorticity $\omega_{0}(r_{0})=10$ in the Eulerian description. According to equation (56), the asymptotic behavior of the stretching rate $\gamma$ is expected to exhibit a highly skewed profile near the boundary as the solution approaches the singularity time. This characteristic is reflected in the spatial distribution observed in the figure. In contrast to $\gamma$’s blowup being focused at the boundary, $\omega$ concentrates its blowup at the center ($r=0$), demonstrating the complementary nature of their singularities. The milder growth rate of $\omega$, evident from the $z$-axis scaling, is consistent with its Lambert-W-dominated asymptotics as predicted by (56).